The problem is: A roller coaster carries 92 people altogether. Some of its cars carry 4 passengers, and the rest carry 6 passengers. There are three fewer 6-passenger cars than 4-passenger cars. How many 4-passenger cars are there?
I started a problem with f=4passenger cars & f-3=6passenger cars, but this doesn't work.

Answer by edjones(8007)

Let f=number of 4 passenger cars, f-3=number of 6 passenger cars.
4f+6(f-3)=92
4f+6f-18=92
10f-18+18=92+18
10f=110
f=11 Ans.
.
Check:
11*4=44
8*6=48
44+48=92
.
Ed